OpenStudy (anonymous):

Can someone help me with this problem? I have to write in point-slope form an equation for the line that goes through each pair of points for (1, 9) and (6, 2)

OpenStudy (esshotwired):

So point-slope form is y-y1=m(x-x1). But first you have to find the slope. The equation to find the slope using two points is y2-y1/x2-x1 or you can do y1-y2/x1-x2. Tell me when you get the slope and I will take you step by step through it.

OpenStudy (anonymous):

So it would be like 1 - 9/6 - 2?

OpenStudy (esshotwired):

Not really. It would be like 9-2/1-6. Do see what I am trying to get at? The y in the first ordered pair and the y in the second ordered pair go on the top, and the x values go on the bottom.

OpenStudy (esshotwired):

But you have to make sure you always line up the ordered pair so it would be wrong if you did 9-2/6-1 because then (6,9) would be an ordered pair but it isn't.

OpenStudy (anonymous):

Ahhh okay.

OpenStudy (esshotwired):

So the slope is 7/-5, but the denominator cannot be negative so you make the 7 negative. m=-7/5

OpenStudy (anonymous):

So denominators are never negative? Just making a note of that. Then what?

OpenStudy (esshotwired):

Then plug in one of the ordered pairs and the slope into the equation y-y1=m(x-x1)

OpenStudy (esshotwired):

This is an example for using the ordered pair (6,2), but you do (1,9). y-2=-7/5(x-6). That is all you do for point-slope form.

OpenStudy (anonymous):

so for mine, using the (1,9) like this? y-9=-7/5(x-1)

OpenStudy (esshotwired):

Yes exactly! :)

OpenStudy (anonymous):

Alright, thanks for the help! :)

OpenStudy (esshotwired):

Your welcome!